OpenStudy (anonymous):

What is the vertex of Y = -x^2 +2x +3 (0, 3) (1, 0) (1,4) (4,1)

OpenStudy (noelgreco):

The x-coordinate of the vertex is found by: \[x=\frac{ -b }{ 2a }\] the front end of the quadratic formula. The y-coordinate is found by simply plugging the x result in the original equation.

OpenStudy (anonymous):

I'm lost, can you explain more please? I got D as my answer

jimthompson5910 (jim_thompson5910):

-x^2 +2x +3 is the same as -1x^2 +2x +3

jimthompson5910 (jim_thompson5910):

-1x^2 +2x +3 is in the form ax^2 + bx + c so that means a = -1 b = 2 c = 3

jimthompson5910 (jim_thompson5910):

Plug a = -1, b = 2 into \[\large x=\frac{ -b }{ 2a }\] to get x = ??

OpenStudy (anonymous):

Alright hold on hun

OpenStudy (anonymous):

I'm lost..

jimthompson5910 (jim_thompson5910):

\[\large x=\frac{ -b }{ 2a }\] \[\large x=\frac{ -2 }{ 2(-1) }\] \[\large x=?\]

OpenStudy (anonymous):

This is like another language hun.. Is there another way to solve it?

jimthompson5910 (jim_thompson5910):

what is 2(-1) equal to

jimthompson5910 (jim_thompson5910):

"2(-1)" means "2 times -1" btw

OpenStudy (anonymous):

Oh! thank you!

jimthompson5910 (jim_thompson5910):

what is x? were you able to find it?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

x = ??

jimthompson5910 (jim_thompson5910):

I'm just making sure you have the right answer

OpenStudy (anonymous):

i got (1,4)

jimthompson5910 (jim_thompson5910):

good here's the graph to confirm

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome